OpenStudy (anonymous):

Which of the following is an equation of the line that passes through the point (1, 1) and has a slope of 2?

OpenStudy (anonymous):

Please post the answers @kayla__

OpenStudy (anonymous):

sorry forgot >_< A. y = 2x – 1 B. y = 2x + 1 C. y = –2x – 1 D. y = –2x + 1

OpenStudy (anonymous):

ugh im doing these 2. they suck dont they! @kayla__

OpenStudy (anonymous):

yes ! im so confused D: @katiebum479

OpenStudy (anonymous):

i am too , i dont mind graphing but they make it all confusing with weird words and stuff. @kayla__

OpenStudy (anonymous):

mmhm. agreed @katiebum479

OpenStudy (anonymous):

B. These are so easy! Just graph them here: https://www.desmos.com/calculator Btw, the slope is the number with 'x'(when it's in the form of y = mx + b)

OpenStudy (anonymous):

ohhhh. thank you ! @greenlegodude57

OpenStudy (anonymous):

You can give medals by clicking 'Best Response'. :P